OpenStudy (anonymous):

48 is 60% of what number?

OpenStudy (paki):

48=60/100*x... solve for "x"....

OpenStudy (anonymous):

what do I multiply it by?

OpenStudy (anonymous):

80*.60=48

OpenStudy (paki):

left "x" alone on one side....and solve then

OpenStudy (anonymous):

\[48=\frac{ 60 }{ 100}*x,x=\frac{ 100*48 }{60 }=?\]

OpenStudy (cwrw238):

you could construct a 'Know Want' table to do these

OpenStudy (cwrw238):